MOA-II Obsevation Log Monthly Summary 2013/03
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 39725-40104 | 373 | 3113 | 200940 |
LMC | 7548-7689 | 134 | 747 | 204500 |
SMC | 1346-1366 | 20 | 40 | 10800 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 527 | 3900 | 416240 (115h00m2240s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
152 | 145 | 322 | 325 | 310 | 50 | 103 | 136 | 289 | 277 | 60 | 66 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
106 | 267 | 63 | 62 | 98 | 98 | 77 | 48 | 45 | 14 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
30 | 33 | 29 | 22 | 22 | 33 | 132 | 127 | 125 | 42 | 33 | 37 |
Field13 | Field14 | Field15 | Field16 |
39 | 43 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
20 | 20 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2013/03/1: / night time = 31680(sec)
2013/03/2: / night time = 31920(sec)
2013/03/3: / night time = 32100(sec)
2013/03/4: / night time = 32340(sec)
2013/03/5: / night time = 32640(sec)
2013/03/6: / night time = 32820(sec)
2013/03/7: / night time = 33060(sec)
2013/03/9: / night time = 33480(sec)
2013/03/10: / night time = 33720(sec)
2013/03/11: / night time = 33900(sec)
2013/03/12: / night time = 34140(sec)
2013/03/13: The filter changer had an error / night time = 34320(sec)
2013/03/14: / night time = 34560(sec)
2013/03/15: / night time = 34740(sec)
2013/03/16: / night time = 31320(sec)
2013/03/17: / night time = 31560(sec)
2013/03/19: / night time = 35580(sec)
2013/03/20: Driver unit error shortly after startup and zero search, resulting in focus change error? V-band filter change was OK / night time = 35760(sec)
2013/03/21: V-band observations sanserror / night time = 35940(sec)
2013/03/22: GB V-band observations went well / night time = 36180(sec)
2013/03/23: / night time = 36360(sec)
2013/03/24: / night time = 36600(sec)
2013/03/25: / night time = 36780(sec)
2013/03/26: / night time = 36960(sec)
2013/03/27: / night time = 37140(sec)
2013/03/28: / night time = 37380(sec)
2013/03/29: / night time = 37560(sec)
2013/03/30: / night time = 37740(sec)
2013/03/31: / night time = 37920(sec)
TOTAL NIGHT TIME: 1006200(sec)
OPERATING RATE: 56.9 %
GB: 32.3 %
LMC: 23.3 %
SMC: 1.2 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month